OpenStudy (anonymous):

1. What is the missing term in the geometric sequence? 3/4, 1/2, 1/3, ? , 4/27, 8/81

OpenStudy (anonymous):

find the ratio

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

All you have to do is find the ratio. You can figure this out by taking any number from the list, and dividing by the number before it. Just take (1/3) / (1/2), and that's the ratio! (1/3) / (1/2) = 2/3 so, 1/3 * 2/3 = 2/9, which is the missing number!

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

not a problem :)